OpenStudy (goformit100):

Use Euclid’s algorithm to find the HCF of 4052 and 12576

OpenStudy (goformit100):

@terenzreignz

terenzreignz (terenzreignz):

I would if I had any idea what Euclid's algorithm is... also, you blocked me?!

OpenStudy (goformit100):

Thankyou sir

OpenStudy (goformit100):

is it correct Step 1 : Since 12576 > 4052, we apply the division lemma to 12576 and 4052, to get 12576 = 4052 × 3 + 420 Step 2 : Since the remainder 420 ≠ 0, we apply the division lemma to 4052 and 420, to get 4052 = 420 × 9 + 272 Step 3 : We consider the new divisor 420 and the new remainder 272, and apply the division lemma to get 420 = 272 × 1 + 148

OpenStudy (mayankdevnani):

Euclid algorithm :